OpenStudy (anonymous):

in fall season league copmetitions, each match pits2 teams agaist each other. if it takes1560 matches for every team to play all the otherteams exctly once, how many teams are playing in the league fall season ? on this data sheet, shoe your idea and thoughts process as you work towrd an answer.

OpenStudy (mathmate):

Same reply as your repost: With 5 teams, each team will play 4 other teams. But this will count each match twice, so the number of matches to play will be 5*4/2. With n teams, the number of matches will be n(n-1)/2. In the present problem, n(n-1)/2=1560. Cross multiply and set up equation: n(n-1)=3120, or n^2-n-3120=0 Solve for n using the quadratic equation.